At 10:30 AM,10:30 AM, detectives discover a dead body in a room and measure its temperature at 27∘C.27∘C. One hour later, the body’s temperature had dropped to 24.8∘C.24.8∘C. Determine how many hours had passed since the death occured (when the body temperature was a normal 37∘C),37∘C), by the time the body was discovered, assuming that the temperature in the room was held constant at 18∘C.at 18∘C.

(Use decimal notation. Give your answer to one decimal place. Round any intermediate calculations, if needed, to no less than three decimal places.)
Answer #1

Arst Given Detective discover a body at 10:30 am in a room whose temperature is 18°c and the body temperature is 27°C, One ho7 (dI = k dt J(18-7) t-tn 118-7)=kt to 7 118 - T1 =ė (k =) T-18= a e-kt-c attoo, T=27 = 27-18 = é o a T-18= gekt at tat ,T= 2